Examine the function, f (x) = (x +1)3 (x − 3)2 for extreme values.


1
Expert's answer
2021-06-09T17:15:45-0400
"f(x)=(x+1)^3(x-3)^2"

Domain: "(-\infin, \infin)"


"f'(x)=((x+1)^3(x-3)^2)'="

"=3(x+1)^2(x-3)^2+2(x-3)(x+1)^3"

"=(x+1)^2(x-3)(3(x-3)+2(x+1))"

"=(x+1)^2(x-3)(5x-7)"

Find the critical number(s)


"f'(x)=0=>(x+1)^2(x-3)(5x-7)=0"

"x_1=-1, x_2=1.4, x_3=3"

Critical numbers: "-1, 1.4, 3"

If "x<-1, f'(x)>0, f(x)" increases.


If "-1<x<1.4, f'(x)>0, f(x)" increases.


If "1.4<x<3, f'(x)<0, f(x)" decreases.


If "x>3, f'(x)>0, f(x)" increases.



"f''(x)=((x+1)^2(x-3)(5x-7))'"

"=2(x+1)(x-3)(5x-7)+(x+1)^2(5x-7)"

"+5(x+1)^2(x-3)"

"=(x+1)(10x^2-14x-30x+42)"

"+(x+1)(5x^2-7x+5x-7)"


"+(x+1)(5x^2-15x+5x-15)"

"=(x+1)(20x^2-56x+20)"

"f''(-1)=0"

"f''(1.4)=-46.08<0"


"f''(3)=128>0"

The function "f(x)" has a local maximum at "x=1.4."


The function "f(x)" has a local minimum at "x=3."


The function "f(x)" has neither maximum nor minimum at "x=-1."